OpenStudy (anonymous):

By what number must you multiply each side of the equation 3/4x=9 to produce the equivalent equation x = 12? 4 4/3 1/4 -3/4

OpenStudy (dtan5457):

To get 3/4x to just x, you multiply by the reciprocal.

OpenStudy (dtan5457):

Which is..?

OpenStudy (anonymous):

the equation is 3/4x=9

OpenStudy (dtan5457):

I know

OpenStudy (anonymous):

Uh. 4/3x=9?

OpenStudy (dtan5457):

3/4x=9 Multiply both sides by 4/3 x=12

OpenStudy (anonymous):

Oh. Okay. Math really confuses me. I'm sorry.

OpenStudy (dtan5457):

It's fine. Good job

OpenStudy (anonymous):

Ok. Thank you. :)